A 4.78 g sample of a laboratory solution contains 1.42 g of acid. What is the concentration of the solution as a mass percentage?
1
Expert's answer
2019-10-21T07:16:20-0400

W=m(acid) /m(sol)*100% =1.42/4.78 * 100% = 29.7%